# A compact neuromorphic system for ultra energy-efficient, on-device robot localization

### v1

@@ Discussion
This work demonstrated a fully neuromorphic visual place recognition system capable of real-time localization, which was successfully deployed on a hexapod robotic platform without any external sensing or compute resources. The LENS spiking neural network model, characterized by its simplicity and compactness, performed accurately across multiple environments and scales, demonstrating its versatility for a range of navigation platforms and tasks. Our system integrated with the SPECK™ chip operated with three orders of magnitude lower energy than conventional von Neumann computing platforms, underscoring its viability for deployment on resource-constrained platforms such as unmanned aerial vehicles  (?), underwater  (?), and locomotive robotics  (?).
@@
Notably, our models were less than 180 KB in size, with the largest containing just 44k parameters. Accurate localization with compact models enables scalability, particularly with deployment on neuromorphic processors with higher memory capability for larger models (Fig. S4). Previous research has shown that accurate place recognition with event cameras can be achieved with as few as 25 pixels  (?). Our system capitalized on this by expanding the feature representation of event pixels, rather than increasing pixel count, to enhance performance (Fig. S4). A limiting factor in this work was the memory capacity of the neurocores for SPECK™, which cap at 64 KB per core. Overcoming this limitation with neuromorphic hardware offering higher memory capabilities could further enhance the performance of large-scale place recognition tasks (Fig. S4). Nonetheless, we have shown comparable performance for place recognition in highly compact spiking network models across multiple environments.
As observed in the comparison between indoor and outdoor localization tasks (Fig. 3), higher dynamic ranges proved challenging for LENS. Future work should focus on dynamically adjusting the biases and parameters of DVS event streams to better handle changes in lighting conditions  (?). Additionally, the turning motions of the hexapod proved challenging for LENS, as rotations triggered widespread activation across the DVS field of view, potentially introducing noise while localizing. This issue was particularly pronounced in outdoor tasks, where the navigation route involved more turns than the indoor path. Continued development of this system will focus on detecting and handling turning motions to for improved localization performance.
@@
Another limitation of our system is the use of static, temporally framed representations generated from event streams without temporal binning for model training. Training our model directly on DVS event streams could help overcome the localization challenges observed in dynamic environments. Developing this would also lend itself for online learning capabilities, where event streams are used directly to train network models in real-time. This would enable the system to adapt to new environments on-the-go, potentially incorporating introspective capabilities that allow the robot to recognize when it is uncertain about its location and learn accordingly  (?, ?, ?, ?).
@@
Event processing
The Locational Encoding with Neuromorphic Systems (LENS) framework leverages spiking neural networks (SNN) on neuromorphic processors for computationally and energy efficient place recognition. We deployed our SNN model on a SynSense Speck2fDevKit, which incorporates the SPECK™ neuromorphic processor with a $128\times 128$ dynamic vision sensor (DVS)  (?) for event driven, real-time localization.
To optimize input size and remove low-activity pixels, we first select an $80\times 80$ region of interest (ROI) in the top-center of the sensor. Events from this ROI were processed using a 2D convolution layer with a kernel size and stride of 8. We set the center weight of the kernel to 1, with all other weights at 0, effectively selecting the center pixel from each convolution to reduce the input to 100 neurons. This selection approach follows a previously established method for event pixel selection  (?).
The input neurons were sparsely connected to a linear feature layer of 200 neurons, with excitatory and inhibitory connection probabilities set at 10% and 50% respectively. The feature layer was fully connected to the output layer, with the number of output neurons corresponding to the number of learnt places (see Materials and Methods: Model training). Pre-processed events were asynchronously collected and passed into our model, where Constant-leak Integrate and Fire (IAF) neurons propagated spikes across the network layers:
@@
Model training
Our network models were trained using temporal representations of event streams, normalized in the range $[0,1]$ and encoded using a time-to-first spike coding scheme  (?). The network was trained based on the temporal dynamics of the spike-timing  (?, ?). The 3-layer architecture (input, feature, and output) used unsupervised learning between the input and feature layers, and a supervised delta learning rule for the feature to output layer.
Unsupervised learning applied spike-timing dependent plasticity (STDP) to update weights $W$, encouraging or pruning connections between neurons. Connections can be positive (excitatory connection) or negative (inhibitory connection):
where $W^{nm}_{ij}$ is the connection weight between neuron $j$ in layer $n$ to neuron $i$ in layer $m$, $\eta_{\text{STDP}}$ is the STDP learning rate, $f_{j}^{n}$ is the target firing rate of neuron $j$ in layer $n$, $\Theta(\cdot)$ is the Heaviside step function, $x^{m}_{i}$ and $x^{n}_{j}$ are the pixel intensity spike values of pre- or post-synaptic neurons respectively, and $t$ is the timestep.
$\begin{gathered}\Delta W^{nm}_{ji}(t)=\eta_{\text{STDP}}(t)/f^{n}_{j}.[x^{m}_{i}(t-1)(x_{\text{force j}}^{n}(t)-x^{n}_{j}(t)],\end{gathered}$ (6)
where $x_{\text{force}}$ is the forced spike in the selected output neuron during training.
The hyperparameters used for our models can be found in Table S1 which were obtained by performing a random hyperparameter sweep search.
@@
Place matching was performed using real-time sequence matching  (?, ?) to enhance the precision of our VPR model on the SPECK™. Spike counts from the output layer were collected in 1-second time bins, then averaged over 4 time bins (4 seconds), representing $\approx$0.5 m of robot movement (average speed of 13.5 $\text{cm}\cdot\text{s}^{-1}$). Once 4 sets of output spike counts were gathered, representing about 2 m of travel, they were compiled into a distance matrix $M_{dist}$, which was convolved with an identity matrix $\mathbb{I}$ to generate a sequence-based distance matrix $M_{seq}$, as previously described  (?):
$\begin{gathered}M_{seq}(i,j)=\mathbb{I}*M_{dist}(i,j)/L.\end{gathered}$ (8)
where $i$ and $j$ are the row and column of the matrix, respectively, and $L$ is the sequence length, which we set to 4 matching the number of output spike counts.
The matched place $\hat{p}$ was identified as the place $i$ corresponding to the neuron with the highest spike count $x_{i}$:
$\begin{gathered}\hat{p}=\operatorname*{arg\,max}_{i}x_{i}.\end{gathered}$ (10)
@@
For power and energy consumption comparisons, we simulated event-driven inference using event frames from the SPECK™ energy experiments (see Materials and Methods: Power and energy measurements). Integrate And Fire neuron spike rates were generated using a time-based rate code:
@@
We used the commercially available JetHexa hexapod (Hiwonder, Shenzen China) customized with a 3D-printed mount to house the Speck2fDevKit (see Fig. 1). Training data was collected by teleoperating the robot for a reference traversal. Real-time localization was performed on the SPECK™ neuromorphic processor with the pre-trained model closely following the reference traversal. Indoor and outdoor datasets represented approximately 40m and 27m of traversal, encoding 75 and 48 unique places, respectively. To ensure accurate ground truth alignment between the mapping and localization phases, experiments were consistently started and ended at the same location.
@@
Intel® CPU power was measured using SoC Watch (Fig. S2), and NVIDIA Jetson Nano power was recorded with the jetson-stats Python package (Fig. S3), with baseline idle power subtracted in both cases. SPECK™ power tracks (IO, RAM, memory, VDDA, and VDDD) were recorded at 20 Hz using the samna package during testing (see Movie 1 and Fig. 2).
@@
We evaluate our system using Recall@N and precision-recall curves, two standard metrics for VPR studies  (?). Recall@N measures the accuracy of the system such that for $N=1$ only the highest match is considered, whereas for $N=5$ the top 5 matches are considered when calculting the accuracy. We used the Sum of Absolute Differences (SAD) method for comparison, which computes pixel-wise similarity across reference and query images and has been used previously for place recognition using event streams  (?, ?). We allow a ground-truth tolerance of $\approx$1m, which is equivalent to $\pm$ 2 reference places.
Code and data availability
To foster future research and development, we have made the code and data for our SNN and its robotic deployment available at:
https://github.com/AdamDHines/LENS
@@ Acknowledgments
This work received funding from an ARC Laureate Fellowship FL210100156 to MM and AUSMURIB000001 associated with ONR MURI grant N00014191-2571. The authors acknowledge continued support from the Queensland University of Technology (QUT) through the Centre for Robotics. We wish to acknowledge the support of the Research Engineering Facility (REF) team at QUT for the provision of engineering support, expertise and research infrastructure in enablement of this project. We also would like to thank Dr Scarlett Raine and Dr Somayeh Hussaini for their valuable feedback during manuscript preparation. Finally, we would like to thank the organisers and participants of the 2022 Lifelong Learning at Scale topic area at the Telluride Neuromorphic Workshop for the insightful discussions and inspiring environment, which helped shape some of the ideas explored in this paper.
@@
Figure S1
Figure S2
Figure S3
Figure S4
Table S1
Movie 1
@@
Figure S2: Power and energy measurements for our Intel® CPU experiment. The baseline power consumption and energy was subtracted using the SoC Watch software from the inferencing measurements for our comparisons to other hardware.
Figure S3: Power and energy measurements for our Jetson Nano experiment. The baseline power consumption and energy measured with the jetson-stats package was subtracted from the inferencing measurements for our comparisons to other hardware.
Figure S4: Effects of model size and sequence length on network performance. (Top) Heatmap of the area under the curve of the Recall@N results from training and testing the Brisbane Event VPR data  (?) for different number of input pixels, feature layer size multiplier (relative to number of pixels), and sequence lengths. Our model with $49\times 49\times 641$ neurons with a sequence length of 10 was the highest performer whilst still being able to be deployed to SPECK™. The best performing model had the architecture of $100\times 200\times 641$ neurons and a sequence length of 10, indicating that model accuracy heavily depends on the number of input events the ratio of the feature layer to the input (Bottom). Performance of the various models relative to sum of absolute differences for the area under the curve.
@@
Values 0.1 0.5 1.0 1.0 128 128

### v2

@@ Discussion
This work demonstrated a fully neuromorphic visual place recognition system capable of real-time localization, which was successfully deployed on a hexapod robotic platform without any external sensing or computational resources. The LENS spiking neural network model, characterized by its simplicity and compactness, performed accurately across multiple environments and scales, demonstrating its versatility for a range of navigation platforms and tasks. Our system integrated with the Speck™ chip operated with 1-2 orders of magnitude lower energy than conventional von Neumann hardware, underscoring its viability for deployment on resource-constrained platforms such as unmanned aerial vehicles  (?), underwater  (?), and locomotive robotics  (?).
@@
Notably, our models were less than 180 KB in size, with the largest containing just 44 K parameters. Accurate localization with our compact model enables portable deployment on neuromorphic processors with higher memory capability for larger models to improve overall performance (Fig. S6). Previous research has shown that accurate place recognition with event cameras can be achieved with as few as 25 pixels  (?). Our system capitalized on this by balancing the number of pixels with the size of the feature layer to achieve the optimal performance (Fig. S6). A limiting factor in this work was the memory capacity of the neurocores for Speck™, which cap at 64 KB per core. Overcoming this limitation with neuromorphic hardware offering higher memory capabilities could further enhance the performance of large-scale place recognition tasks (Fig. S6). Nonetheless, we have shown comparable performance for place recognition in highly compact spiking network models across multiple environments.
Our neuromorphic place recognition system employs real event streams directly for localization in a unique way, establishing a new methodology for event-driven place recognition. This work highlights the potential for substantial efficiency gains when using event-based DVS over conventional cameras, especially in scenarios where energy-efficient processing is critical. Unlike frame based cameras, which operate continuously, the event-driven DVS captures only essential changes, reducing data redundancy and energy consumption. Our experiments encompass both a conventional benchmark dataset widely used in the robotics community  (?) and a more challenging legged robot navigation task, ensuring that the localization accuracy of our method is evaluated across a diverse range of environments and baseline performance levels. The accuracy achieved by LENS demonstrates its effectiveness within broader systems such as Simultaneous Localization and Mapping (SLAM), where geometric constraints and filtering methods can mitigate occasional inaccuracies  (?, ?, ?). By utilizing a more stable platform, such as a four wheeled robot, our network is able to achieve high precision comparable to a benchmark technique. This demonstrates that our compact spiking network models perform comparably to conventional non-event-driven methods, highlighting the potential for further improvements with future developments in neuromorphic hardware and algorithmic enhancements.
As observed in the similarity matrices for the localization tasks (Fig. 5), introducing many turns into shorter trajectories proved challenging for LENS in the outdoor hexapod experiment. Dynamically adjusting these biases and parameters of DVS event streams to better handle these sudden motions could enhance accuracy across challenging environments  (?). Continued development of this system will focus on detecting and handling turning motions to improve localization performance.
@@
Another limitation of our system is the use of static, temporal representations of places generated from event streams using time-window binning for model training (Fig. 2). Training our model directly on DVS event streams could help overcome the localization challenges observed in dynamic environments. Developing this would also facilitate online learning capabilities, where event streams are used directly to train network models in real-time. This would enable the system to adapt to new environments on-the-go, potentially incorporating introspective capabilities that allow the robot to recognize when it is uncertain about its location and learn accordingly  (?, ?, ?, ?). Additionally, using more information from the DVS sensor such as including convolutional layers and self-supervised learning could be useful to extract more features for enhanced performance  (?).
Integration of our neuromorphic system with other methods for navigation could be used to enhance and improve robotic localization performance. Dupeyroux et al. developed a bio-inspired path integration system for a hexapod robotic platform that would serve as a complementary method for VPR  (?). Additionally, other insect inspired systems employ methods such as goal approaching and collision avoidance which allows for autonomous navigation in novel environments, typically for homing tasks  (?, ?, ?). Integration of these mixed-model paradigms present themselves as a promising method to be used in more comprehensive navigation systems such as SLAM, of which there are many neuroscience inspired systems available  (?, ?, ?, ?).
In conclusion, we present a fully end-to-end neuromorphic VPR system that performs accurate, low-power localization on robotic platforms. Our work has the potential to enable robots in the field to navigate further and conduct longer missions by reducing the computational and energy demands of current vision systems  (?). An important aspect of this work was to ensure that neuromorphic localization could be achieved on multiple robotic platforms, including for resource-constrained scenarios. Our approach moves VPR toward efficient, versatile neuromorphic localization, unlocking new possibilities for autonomous robotic navigation at the edge.
@@
Model training
Our network models were trained on CPU hardware or a Jetson Nano using static temporal representations of event streams with pixel intensity values relating to the number of events collected over a one second time-window, subsequently normalized in the range $[0,1]$ (see Fig. 2). The network was trained using spike timing from an abstracted spiking method, where the pixel intensity value represents temporal activity  (?, ?, ?). The networks three layer architecture (input, feature, and output) used unsupervised learning between the input and feature layers, and a supervised delta learning rule for the feature to output layer.
Unsupervised learning applied spike-timing dependent plasticity (STDP) to update weights $W$, encouraging or pruning connections between neurons. Initial weights and connections were randomly generated from predetermined probabilities, with connections either being positive (excitatory connection) or negative (inhibitory connection). From the input to feature layer there was a $35\%$ and $75\%$ excitatory and inhibitory connection probability, respectively. The feature to output layer was fully connected with an equal distribution of excitatory and inhibitory connections. As spikes were propagated through the network, the layer weights were updated in the following way:
where $W^{nm}_{ij}$ is the connection weight between neuron $j$ in layer $n$ to neuron $i$ in layer $m$, $\eta_{\text{STDP}}$ is the STDP learning rate, $f_{j}^{n}$ is the target firing rate of neuron $j$ in layer $n$, $\Theta(\cdot)$ is the Heaviside step function, $x^{m}_{i}$ and $x^{n}_{j}$ are the neuron states $x$ in the network layer $i$ and $j$ with connections $m$ and $n$, and $t$ is the timestep. During weight updates throughout training, we did not specify any rule that says weights must remain excitatory or inhibitory, however we did not observe weights switching sign. That is, initially seeded connection types remained robust throughout training.
$\begin{gathered}\Delta W^{nm}_{ji}(t)=\eta_{\text{STDP}}(t)/f^{n}_{j}.[x^{m}_{i}(t-1)(x_{\text{force j}}^{n}(t)-x^{n}_{j}(t))],\end{gathered}$ (4)
where $x_{\text{force}}$ was the forced spike in the selected output neuron during training.
The hyperparameters used for our models can be found in Table S1 which were obtained by performing a hyperparameter sweep.
Deployment time
The Locational Encoding with Neuromorphic Systems (LENS) framework leveraged spiking neural networks (SNN) on neuromorphic processors for computationally and energy efficient place recognition. We deployed our pre-trained SNN model on a SynSense Speck2fDevKit, which incorporates the Speck™ neuromorphic processor with a $128\times 128$ dynamic vision sensor (DVS)  (?) for event driven, real-time localization. This was in contrast to the training methodology where event frames were generated by counting the number of events over a specified time-window. Our localization system used truly asynchronous event streams to continually bias and activate neurons as the robot navigated through the environment.
To optimize input size and remove low-activity pixels, we first selected an $80\times 80$ region of interest (ROI) in the top-center of the sensor. Events from this ROI were processed using a 2D convolution layer with a kernel size and stride of 8, which acted as the first processing layer of our SNN prior to the pre-trained input layer and was deployed on chip. We set the center weight of the kernel to 1, with all other weights at 0, effectively selecting the center pixel from each convolution to reduce the input to 100 neurons. This selection approach followed a previously established method for event pixel selection  (?).
The input neurons were sparsely connected to a linear feature layer of 200 neurons, with excitatory and inhibitory connection probabilities set at $35\%$ and $75\%$ respectively. The feature layer was fully connected to the output layer, with the number of output neurons corresponding to the number of learned places (see Materials and Methods: Model training). Events were asynchronously collected and passed into our model, where Constant-leak Integrate and Fire (IAF) neurons propagated spikes across the network layers:
@@
Place matching was performed using real-time sequence matching  (?, ?) to enhance the precision of our VPR model on the Speck™. Spike counts from the output layer were collected in one second time bins, then averaged over four time bins (four seconds), representing $\approx$0.5 m of robot movement (average speed of 13.5 $\text{cm}\cdot\text{s}^{-1}$). All place and sequence matching was performed off-chip on a Jetson Nano during on-robot deployment or a CPU for simulated experiments. Once four sets of output spike counts were gathered, representing about 2 m of travel, they were compiled into a distance matrix $M_{dist}$ (Fig. 1D), which was convolved (*) with an identity matrix $\mathbb{I}_{L}$ to generate a sequence-based distance matrix $M_{seq}$, as previously described  (?):
$\begin{gathered}M_{seq}(i,j)=\mathbb{I}_{L}*M_{dist}(i,j)/L,\end{gathered}$ (8)
where $i$ and $j$ are the row and column of the matrix, respectively, and $L$ is the sequence length, which we set to 4 matching the number of output spike counts. For the Brisbane Event Dataset, we set the sequence length to 30. The choice of sequence length matters for both computational efficiency and evaluation performance, with shorter lengths favoring the former and longer lengths benefiting the latter (Fig. S7).
After generating the sequence-based distance matrix $M_{\text{seq}}$, we performed place matching on the processed spike outputs. The matched place $\hat{p}$ was identified as the place $i$ corresponding to the neuron with the highest spike count $x_{i}$ for each query in the $M_{\text{seq}}$:
$\begin{gathered}\hat{p}_{k}=\arg\max_{i}M_{\text{seq}}(k,i)\quad\text{for each}\quad k\in{1,2,3,4}.\end{gathered}$ (10)
where $\hat{p}_{k}$ is the matched place for the $k$-th sequence, $i$ ranged over all database reference places, $M_{\text{seq}}(k,i)$ represents the sequence-based distance between the $k$-th spike count set and the $i$-th reference place.
@@
For power and energy consumption comparisons using LENS on neuromorphic and von Neumann hardware, we simulated event-driven localization using event frames from the Speck™ energy experiments (see Materials and Methods: Power and energy measurements). Integrate And Fire neuron spike rates were generated using a time-based rate code:
@@
We used the commercially available JetHexa hexapod (Hiwonder, Shenzen China) customized with a 3D-printed mount to house the Speck2fDevKit (see Fig. 1). We also simulated a four wheeled robotic deployment by evaluating our system on the Brisbane Event Dataset  (?), to highlight that our energy-efficiency performance is robotic platform agnostic. Training data was collected by teleoperating the robot for a reference traversal. Real-time localization was performed on the Speck™ neuromorphic processor with the pre-trained model closely following the reference traversal. Indoor and outdoor datasets represented approximately 40 m and 27 m of traversal, encoding 75 and 48 unique places, respectively. To ensure accurate ground truth alignment between the mapping and localization phases, experiments were consistently started and ended at the same location.
@@
Intel® CPU power was measured using SoC Watch, and NVIDIA Jetson Nano power was recorded with the jetson-stats Python package, with baseline idle power subtracted in both cases to only measure energy consumed for the VPR algorithms. Speck™ power tracks (IO, RAM, memory, VDDA, and VDDD) were recorded at 20 Hz using SynSense’s samna package during testing (see Supplementary Movie 1 and Fig. 3).
@@
We evaluated our system using Recall@N and precision-recall curves, two standard metrics for VPR studies  (?). All statistical analysis was performed on CPU hardware. Our implementation of precision and recall matched methods used in other place recognition systems  (?) (Fig. S2). Recall@N measured the accuracy of the system such that for $N=1$ only the highest match was considered across all reference places, whereas for $N=5$ the top 5 matches were considered when calculating the accuracy, up to a maximum of 25 potential places. Precision-recall curves were utilized to assess performance by illustrating the relationship between precision and recall across different threshold levels for positive matches, calculated as follows:
$\begin{gathered}Precision=\frac{TP}{TP+FP}\hskip 10.00002ptRecall=\frac{TP}{GTP}\end{gathered}$ (14)
where $TP$ is the number of true positives, $FP$ is the number of false positives, and $GTP$ is the number of ground truth positives  (?). Initially precision and recall is undefined due to there being no predictions, however we initialize precision at 1 and recall at 0 for visualization purposes.
We used the Sum of Absolute Differences (SAD) method for comparison, which computed pixel-wise similarity across reference and query images and has been used previously for place recognition using event streams  (?, ?). Pseudo ground-truth data was collected following  (?) using GPS correspondence between reference and query images. Pseudo-GPS and manually confirmed ground-truth showed high correspondence, with the majority of places being ± 1 place label away (Fig. S3). Further analysis of the accuracy of the pseudo ground-truth and applied tolerances were confirmed by near identical results (Fig. S4). We allowed a ground-truth tolerance of $\approx$1 m, which is equivalent to $\pm$ 2 reference places, and a tolerance of $\approx$42 m for the Brisbane Event VPR dataset, which is equivalent to $\pm$ 3 reference places. A visualization of the ground truth, TP, and FP matches for each similarity matrix can be found in Fig. S5.
Data and code availability
All data needed to support the conclusions of this manuscript are included in the main text, Supplementary Materials, and are also available alongside the code at https://doi.org/10.5281/zenodo.15392412.
@@ Acknowledgments
The authors acknowledge continued support from the Queensland University of Technology (QUT) through the Centre for Robotics. We wish to acknowledge the support of the Research Engineering Facility (REF) team at QUT for the provision of engineering support, expertise and research infrastructure in enablement of this project. We also would like to thank Dr Scarlett Raine and Dr Somayeh Hussaini for their valuable feedback during initial manuscript preparation. Finally, we would like to thank the organisers and participants of the 2022 Lifelong Learning at Scale topic area at the Telluride Neuromorphic Workshop for the insightful discussions and inspiring environment, which helped shape some of the ideas explored in this paper.
Funding:
This work received funding from an ARC Laureate Fellowship FL210100156 to MM, AUSMURIB000001 associated with ONR MURI grant N00014191-2571 to MM and TF, and an ARC Discovery Early Career Researcher Award DE240100149 to TF.
Author contributions:
A.D.H & T.F. conceptualized and designed the experiments; A.D.H performed and analyzed all experiments; T.F. & M.M. provided general input into the research; T.F. & M.M. provided funding for the project; A.D.H wrote the paper; A.D.H., M.M. & T.F. edited the paper.
Competing interests:
The authors declare no competing interests.
Data and materials availability:
All data and code used in this study is available at https://doi.org/10.5281/zenodo.15392412.
@@
Supplementary Figures S1-S8
Supplementary Tables S1-S2
Supplementary Movie S1
Supplementary Materials for
A compact neuromorphic system for ultra-energy-efficient, on-device robot localization
Adam D Hines1∗, Michael Milford1, Tobias Fischer1
1QUT Centre for Robotics, Queensland University of Technology, 2 George St, Brisbane, QLD 4001, Australia
Corresponding author. Email: adam.hines@qut.edu.au
This PDF file includes:
Supplementary Figures S1-S8
Supplementary Tables S1-S2
Other Supplementary Materials for this manuscript:
Supplementary Movie S1
Figure S1: Power and energy measurements for running LENS on von Neumann hardware. The baseline power consumption and energy was subtracted from the localization measurements for our comparisons to other hardware.
Figure S2: Comparison of Precision-Recall calculation without ground truth tolerance. PR curves for the Brisbane Event-VPR dataset generated by our pipeline (blue) align closely with those generated using the scikit-learn precision_recall_curve function (orange) employed by Yu et al.  (?). The minor deviation is due to discretization artefacts, as our method uses 100 points while the approach taken by Yu et al. uses 398 points to compute the curves.
Figure S3: Comparison of GPS and human-labeled ground truth for queries and their matching references. (Left) Shows the query images with an inset of their place index for the Brisbane Event-VPR dataset. (Middle) The matching reference place using the GPS pseudo ground truth and the (Right) actual match confirmed by manual ground truth. In the cases where Difference = 0, the GPS pseudo ground truth matches exactly to the manually labeled correct image. In cases where Difference=+/- 1, the GPS match was offset by one image of the true match, which is equivalent to approximately 14 m. The purpose of the ground truth tolerances is to allow such neighboring images to count as correct matches, accounting for such slight misalignments.
Figure S4: Precision-recall evaluation comparing GPS pseudo ground truth with manually aligned ground truth. For the Brisbane Event-VPR dataset, we compare the GPS pseudo ground truth (GT) with the manually annotated ground truth using tolerances of 1 and 3 places representing approximately 14 m and 42 m, respectively. Both performed very similarly with a Recall@1 of 0.86 and 0.87 for the manual and GPS pseudo ground truth, respectively. The slight improvement in performance observed with the pseudo ground truth can be attributed to the higher tolerance threshold of 42 m capturing 7 additional matches, which extends beyond the 14 m limit used for manual ground truth correction.
Figure S5: Distribution of true and false positives for similarity matrices. Similarity matrices for (left) LENS and (right) SAD for the 3 datasets evaluated from top to bottom - Brisbane Event Dataset, hexapod indoor, and hexapod outdoor. Matrices show the ground truth used (white dots) with true positives (green circles) and false negatives (red crosses).
Figure S6: Effects of model size and sequence length on network performance. (Top) Heatmap of the area under the curve of the Recall@N results from training and testing the Brisbane Event VPR data  (?) for different number of input pixels, feature layer size multiplier (relative to number of pixels), and sequence lengths. Our model with $49\times 49\times 641$ neurons with a sequence length of 10 was the highest performer whilst still being able to be deployed to Speck™. The best performing model had the architecture of $100\times 200\times 641$ neurons and a sequence length of 10, indicating that model accuracy heavily depends on the number of input events the ratio of the feature layer to the input. (Bottom) Performance of the various models relative to sum of absolute differences for the area under the curve. Smaller number of input pixels with larger feature layer sizes can perform as well as or better than a larger number of input pixels with the same feature layer size. The important consideration is that increasing the feature layer size greatly increases the model size, which may limit deployment on memory constrained hardware.
Figure S7: Effect of sequence length on Recall@1 performance. An ablation study was conducted on the Brisbane Event-VPR dataset observing the effects of sequence length for Recall@1 performance between our system (LENS) and the Sum-of-Absolute-Differences (SAD) baseline method. We used the manually generated ground truth with a tolerance of +/- 1 place (approximately 14 m) for this investigation.
@@
Values 0.35 0.75 1.0 1.0 64 128
Table S2: Energy consumption for SAD with fewer pixels
Pixel count 100 64 1
Energy CPU (mJ) 61,427.0 46943.18 25698.37
Energy Jetson (mJ) 2,968.0 531.81 474.88
Caption for Supplementary Movie S1.
LENS display during localization. The left side shows the main display of LENS during online localization deployed on a robotic platform. The Visualizer shows the field of view of the on-board dynamic vision sensor with the power being actively consumed dynamically plotted underneath. The right side shows the output of the LENS matching system which returns the place matching result after collecting output spikes from the model over the specific time window.